What is the force on the charge located at x = +8.
Answer #1

0 be bbg b+

Problem 7)

The expression for the electric force

F = k q1q2 / r2

The net force about charge -aq

F = F12 – F32

calculate the distance

F12 = k q1 q2 / r122

r12 = 8 -3 = 5 cm = 5 10-2 m

r32 = 11-8 = 3 cm = 3 10-2 m

calculate the force

Data

q = 5 μC = 5 10-6 C

a = 7.50

F12 = k q1 q2 / r122

F12 = 8.987 109 5 10-6 (7.50 5 10-6 )/ (5 10-2)2

F12 = 6.74 102 N

F32 = k q3 q2 / r322

F32 = 8.987 109 5 10-6 (7.50 5 10-6 )/ (3 10-2)2

F32 = 1.87 103 N

The net force is

F = 6.74 102 N - 1.87 103 N

F = -1198 N = -1.20 N

The negative sign indicates that the force is directed to the left

Problem 8

The expression for the electric field is

E = k q / r2

r = 25 cm = 0.25 m

Eb = 8.987 109 10 10-6 /0.252

Eb = 1.44 106 N/C

Ec= 8.987 109 5 10-6 / 0.252

Ec = 7.19 105 N/C

These field are in the line of the triangle we use trigonometry to find components

Cos 30 = Eby/Eb

Eby = Eb Cos 30

Eby = 1.44 106 Cos 30

Eby = 1.247 106 N/C

Sin 30 = Ebx /Eb

Ebx = Eb Sin 30

Ebx = 1.44 106 Sin 30

Ebx = 0.72 106 N/C

Ecy = Ec Cos 30

Ecy = 6.227 105 N/C

Ecx = Ec Sin 30

Ecx = 3.595 105 N/C

The net field is the sum of each axis

Ex = Ebx + Ecx

Ex = 0.72 106 + 3.595 105

Ex =10.795 105 N/C

Ey = Eby – Ecy

Ey = 12.47 105 - 6.227 105

Ey = 6.243 105 N/C

E = sqrt (Ex2 + Ey2)

E = 105 sqrt (10.7952 + 6.2432 )

E = 12.47 105 N/C

Direction

Tan θ = Ey/Ex

Tan θ = 6.243/10.795

θ = 30

Part b)

F = q E

F = 1.5 10-9 12.47 105

F = 1.87 10-3 N

As the charge is positive the direction does not change

θ = 30